OpenStudy (anonymous):

How do I solve by factoring: 2y^2 + 5y = -2?

OpenStudy (anonymous):

I got y^2 + 5/2y + 1 =0 now what?

OpenStudy (anonymous):

you could use the quadratic formula but that is time consuming ...

OpenStudy (anonymous):

here's how I solve it : 2y^+5y+2 =0

OpenStudy (anonymous):

I have to use factoring..

OpenStudy (anonymous):

Take a and c and multiply them : 2*2 = 4

OpenStudy (anonymous):

so I don't divide by 2 to get y^2?

OpenStudy (anonymous):

b is the sum ... look for factors of 4 positive and negative that add up to 5...

OpenStudy (anonymous):

you don't have to ...

OpenStudy (anonymous):

factors of 4 adding up to 5 are 4 and 1 ...right?

OpenStudy (anonymous):

Now... substitute b with the two terms we found (4 and 1 )

OpenStudy (anonymous):

2y^2 +4y +1y +2 =0 same thing right?

OpenStudy (anonymous):

now factor by grouping ...can you do that?

OpenStudy (anonymous):

2y(y+2)+1(y+2) =0 and factor the GCF -> (y+2)(2y+1)=0

OpenStudy (anonymous):

yeah I can thank you!

OpenStudy (anonymous):

You're welcome ...this method works for all trinomials with leading coefficient ≠1

OpenStudy (anonymous):

btw you have to divide by 2 when you have to solve by Completing the Square .

OpenStudy (anonymous):

factor the GCF -> (y+2)(2y+1)=0? What do you mean?

OpenStudy (anonymous):

ohh, okay.